A rod 3 m long is found to have expanded 0.091 cm in length after a temperature rise of 60 deg C. What is the temperature coefficient for the material of the rod?


1
Expert's answer
2021-04-30T11:26:01-0400

Write the equation that determines the amount of thermal expansion:


"L_e=L_0(1+\alpha T),\\\space\\
\alpha=\frac 1T\bigg(\frac{L_e}{L_0}-1\bigg)=5.06·10^{-6}1/°\text{C}"